Giving that mean=34, median= 32 and standard deviation as 12. Calculate the Pearson's coefficient of skewness
0.1
0.4
0.5
1.6
1
Expert's answer
2015-02-27T10:28:00-0500
Information on mode is absent to compute Pearson's first skewness coefficient equal to 3*(mean-mode)/standard deviation. Pearson's second skewness coefficient is 3*(mean-median)/standard deviation=3*(34-32)/12=3*2/12=6/12=0.5.